OpenStudy (isabel☺):

solve y^2 dy = x(x dy - y dx)

OpenStudy (isabel☺):

(DE)

OpenStudy (isabel☺):

@DLS

OpenStudy (isabel☺):

@paki

hartnn (hartnn):

homogeneous DE, so just plug in \(\Large y = vx \\ \Large dy = vdx +x dv \) simplify and it will get converted into easily separable DE, try it :)

OpenStudy (isabel☺):

when should we use y=vx and x=vy?

hartnn (hartnn):

non-homogeneous DE both works.

OpenStudy (isabel☺):

we can use them for non homo and for homo. DE?

hartnn (hartnn):

no, i meant you can use both y=vx and x=v but for only homogeneous eq. For non-homogeneous eq, different metod are there

OpenStudy (isabel☺):

i got \[\ln \frac{ x ^{2} }{ y } = C + \frac{ x ^{2} }{ 2y ^{2} }\] ???

hartnn (hartnn):

i think ln x is getting cancelled.... can you check your work again ?? or show your attempt, i will help you spot the error :)

OpenStudy (dls):

y^2 dy = x(x dy - y dx) dy(y^2-x^2)=-xy dx \[\LARGE \frac{dy}{dx}=\frac{xy}{x^2-y^2}\] put y=vx \[\LARGE v+ x \frac{dv}{dx}=\frac{v}{1-v^2}\] \[\LARGE x \frac{dv}{dx}=\frac{v}{1-v^2} - \frac{v(1-v^2)}{1-v^2}\] \[\LARGE x \frac{dv}{dx}=\frac{v^2}{1-v^2}\] \[\LARGE \frac{1-v^2 }{v^2} dv = \frac{dx}{x}\] Now integrate both sides and compare the starting steps I did with urs.

hartnn (hartnn):

why u do dis dls ;-;

OpenStudy (dls):

i didn't write the whole solution hartnn,just the starting steps

OpenStudy (isabel☺):

uhmm mine is